How do you put this in point slope form? (-2,5) (9,-6)
Answer by richwmiller(17219)

first find the slope
(y-y)/(x-x)=m the slope
-6-5/9+2
-11/11=-1=m
y=mx+b
now plugin one x and one y
and solve for b
5=-1*-2+b
5=3+b
2=b
y=-1x+2
you can plug in the x's and y's to check